[A freight train leaves town at 40 mph. A second train leaves town 3 hours later at 56 mph. How many hours after the first train leaves will the second overtake the first train?]
Found 2 solutions by rfer, TimothyLamb:
Answer by rfer(16322)

Train A has a 120 mile lead.
Overtake speed is 16mph
120/16=7.5 hrs

Answer by TimothyLamb(4379)

s = d/t
d = st
t = d/s
---
distance between trains when 2nd train leaves station:
d = st
d = 40*3
d = 120 miles
---
speed of 2nd train relative to 1st train:
s = 56 - 40
s = 16 mph
---
time for trains to meet:
t = d/s
t = 120/16
---
answer:
t = 7.5 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php